OpenStudy (anonymous):

a supermarket display of cans has 20 cans in the bottom row and 1 can on the top row . Each row , except the bottom row , has 1 can less than the row below it . How many cans are on display?

OpenStudy (anonymous):

@amistre64 help sir with this last question

OpenStudy (anonymous):

who knows this?

OpenStudy (anonymous):

20 + 19 + 18 + 17 + 16 + 15 + 14 + 13 + 12 + 11 + 10 + 9 + 8 + 7 + 6 + 5 + 4 + 3 + 2 + 1 = 210 There are 20 rows with a total of 210 cans. There is probably an easier way to do it, I just don't know the equation for it.

OpenStudy (anonymous):

oh tnx kellie :D

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

u could have used this formula =n(n+!)/2=20(21)/2+210 ( which applies only for natural number from 1 with difference 1) instead of adding it one by one

OpenStudy (anonymous):

that would have been easier...lol :)

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

tnx im done :D